Find the function (do not submit a graph) that is finally graphed after the following transformations are applied to the given function. Please be sure to show all of your work.
f(x)= 1/x
Shift left 1 unit, shift up 3 units, reflect about the y-axis.

Answer by stanbon(75887)

Find the function (do not submit a graph) that is finally graphed after the following transformations are applied to the given function. Please be sure to show all of your work.
f(x)= 1/x
Shift left 1 unit
---
a(x) = 1/(x+1)
-------
shift up 3 units
b(x) = 1/(x+1) + 3
-----------------------------
reflect about the y-axis.
c(x) = -1/(x+1)-3
==========================
Cheers,
Stan H.